OpenStudy (anonymous):

i'll fan and medal you Solve this system of equations using any method. 3x – y = 4 x + 2y = 13

OpenStudy (anonymous):

@Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

I actually have no idea how to do this. o.O I should know, but I completely forgot how to solve this type of problem. Dx Maaaaaaaybe @Novalynne can help. c:

OpenStudy (anonymous):

@esshotwired can you help or

OpenStudy (esshotwired):

Yes, I can help ;)

OpenStudy (anonymous):

ess pls

OpenStudy (esshotwired):

>_> fine gimme a second

OpenStudy (dan815):

hey can you rearrange one of the equations in terms of x or y?

OpenStudy (esshotwired):

So first you want to try and make either the x or y values the same but with opposite signs. For this one, it would be pretty easy to do the y value, so multiply the top equation by 2. (3x-y=4)2 6x-2y=8 Now take that new equation and the second original equation and add them together. x+2y=13 6x-2y=4 --------- 7x = 17 Does that make sense?

OpenStudy (anonymous):

Did you ever solve for x and y?